x-y=2+2y

x-y=2+2y
 

Grade:8

2 Answers

Jonathan Jackson
9 years ago
I don’t know  exactly what do you need but 
x-y=2+2y
 x-2=3y
x-3y=2 this is what you can do 
you can’t do it futher
Jonathan Jackson
9 years ago
I don’t know  exactly what do you need but 
x-y=2+2y
 x-2=3y
x-3y=2 this is what you can do 
you can’t do it futher